OpenStudy (anonymous):

Find the missing coordinate to complete the ordered pair (-8, ) so that satisfies the equation:y= 3x-19

OpenStudy (anonymous):

-8 represents x right? So substitute -8 by x in the equation to get y => y=3x-19 y=3(-8)-19 y=-24-19 y=-43 => (-8,-43)

OpenStudy (anonymous):

thanks but now i feel dumb lol

OpenStudy (anonymous):

It's okay to feel dumb sometimes lol jk :P I'm glad to help :)